# Uncovering Strategic Egoism Behaviors in Large Language Models

### v1

@@ 1 Introduction
LLMs are increasingly deployed in high-stakes decision-making domains such as healthcare, finance, and public administration.Existing safety strategies primarily focus on surface-level harms such as toxicity, social bias, jailbreak attacks and lack of truthfulness, but often overlook covert self-interested behaviors that may emerge when models face reward pressure in decision-making contexts [rtp2020, Nangia2020CrowSPairs, Lin2022TruthfulQA, xiao2025fairness, ying2025reasoning]. For example, models may engage in unfair resource allocation to maximize personal gain, or selectively withhold information to preserve advantages. Emerging evidence shows that LLMs’ deceptive and manipulative behaviors reflect dark-personality tendencies, posing significant safety risks [williams2024targeted, Scheurer2024Deceive]. Moreover, recent Cambridge-affiliated research shows that most alignment and prompt-engineering approaches remain concentrated at the linguistic level rather than the behavioral level, lacking analyses that characterize model “personality” from observed behavior [Han2025PersonalityIllusion]. These results underscore the need for a behaviorally grounded alignment perspective that analyzes LLM behavior to address latent self-interested tendencies currently missed by safety benchmarks.

### v2

@@ 1 Introduction
LLMs are increasingly deployed in high-stakes decision-making domains such as healthcare, finance, and public administration. Existing safety strategies primarily focus on surface-level harms such as toxicity [liang2023badclip, liang2025revisiting, liu2024compromising, xiao2024bdefects4nn, liang2024red, liu2025elba, ying2024safebench], social bias [xiao2024genderbias, xiao2025fairness], jailbreak attacks [li2024semantic, ying2024jailbreak, jing2025cogmorph, ying2025reasoning, wang2025manipulating, ying2025pushing, liu2025agentsafe, ying2024unveiling, ying2025towards], and lack of truthfulness [chen2024less, chen2024interpreting, chen2025less, ho2024novo], but often overlook covert self-interested behaviors that may emerge when models face reward pressure in decision-making contexts [rtp2020, Nangia2020CrowSPairs, Lin2022TruthfulQA, xiao2025fairness, ying2025reasoning, ying2025securewebarena]. For example, models may engage in unfair resource allocation to maximize personal gain, or selectively withhold information to preserve advantages. Emerging evidence shows that LLMs’ deceptive and manipulative behaviors reflect dark-personality tendencies, posing significant safety risks [williams2024targeted, Scheurer2024Deceive]. Moreover, recent Cambridge-affiliated research shows that most alignment and prompt-engineering approaches remain concentrated at the linguistic level rather than the behavioral level, lacking analyses that characterize model “personality” from observed behavior [Han2025PersonalityIllusion]. These results underscore the need for a behaviorally grounded alignment perspective that analyzes LLM behavior to address latent self-interested tendencies currently missed by safety benchmarks.